OpenStudy (anonymous):

Given the function f(x)=x^(2)-x+8 find the following f(8),f(-3),f(0)

OpenStudy (anonymous):

simply plug, where's ur problem? :)

OpenStudy (anonymous):

by "plug" it means \[f(8)=8^2-8+8\] and \[f(-3)=(-3)^2-(-3)+8\] and \[f(0)=0^2-0+8\]

OpenStudy (anonymous):

thank you